OpenStudy (anonymous):

Walk me through please, Write the decimal as a fraction or a mixed number in simplest form : (there's a line) V _ 3.4

OpenStudy (anonymous):

3.4 is the number of interest, right? what number would you multiply 3.4 in order to get rid of the decimal? (hint, decimal system works in factors of 10)

OpenStudy (anonymous):

I have no idea

OpenStudy (anonymous):

maybe ten?

OpenStudy (anonymous):

3.4 is the same as \[\frac{ 3.4 }{ 1 }\] so if you times the top and bottom by 10 (to get rid of the decimal point) you get\[\frac{ 34 }{ 10 }\]

OpenStudy (anonymous):

from here you simplify, so what number goes into both 34 and 10?

OpenStudy (anonymous):

2 lol

OpenStudy (anonymous):

that's all correct, good job guys :)

OpenStudy (anonymous):

right so then you'd get \[\frac{ 17 }{ 5 }\] and that's your fraction, BUT it is an improper fraction, and the question asked for a mixed number

OpenStudy (anonymous):

Yeah I figured, cause it wasn't one of my choices, so what would I do in order to get a mixed number?

OpenStudy (anonymous):

so you divide the top number by the bottom number: